OpenStudy (anonymous):

c2-10c=0? what are the solutions?

OpenStudy (anonymous):

+or-sqrt(10)

jhonyy9 (jhonyy9):

factoriz out c ,so than what will get ?

Parth (parthkohli):

Factor. \(\Large \color{MidnightBlue}{\Rightarrow c(c - 10) = 0 }\) By the property 0x = 0 Either c = 0 -OR- c - 10 = 0 c = 10 The correct solution set of x is (0,10)

Parth (parthkohli):

of c*